OpenStudy (anonymous):

Which expressions show how 6 x 45 can be rewritten using the distributive property? A. 6 x 40 + 6 x 5 B. 6 x 40 − 6 x 5 C. 6 x 50 – 6 x 5 D. 6 x 50 + 5

OpenStudy (anonymous):

choose all answers that are correct

OpenStudy (cwrw238):

6 * 45 = 6(40 + 5) thats one (A) can you find the other one?

OpenStudy (anonymous):

@cwrw238 um no help please :(

OpenStudy (anonymous):

@cwrw238

OpenStudy (anonymous):

@cwrw238 i found it it's c right?

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

6 (50 -5) = 6 * 45

OpenStudy (cwrw238):

6 * 45 = 270 6*50 - 6*5 = 300 - 30 = 270